# ![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png) CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces | Adrian Xiao Bin Yong, and Elif Ertekin |

### v1

@@ 2.2 Dataset generation
To reduce the time required for DFT relaxations, we trained from scratch a machine learning interatomic potential, M3GNet 48, for the LSC-LCO interfaces, and used to it to pre-relax structures prior to DFT relaxation. A total of 15,484 training structures consisting of optimized structures and intermediate ionic steps of the DFT relaxations were used to train the M3GNet interatomic potential. The M3GNet model achieved low test set mean absolute errors (MAEs) of 2.70 mev/atom, 20.9 meV/Å, and 0.0146 GPa for energy, force, and stress respectively. Machine learning interatomic potentials with similar (or higher) MAEs showed good performance in relaxations and molecular dynamics simulations when applied to other Li-ion conductors 48, 49. We then relaxed randomly generated interface structures using the M3GNet interatomic potential, followed by subsequent DFT relaxations. The interface energies of the relaxed structures were distributed across a wide range (approximately 1.4 J/m) as shown in ESI Fig. S1. We defined structures with interface energies no higher than 0.4 J/m relative to the lowest energy structure to be low-interface-energy structures. For reference, the observation frequency of a grain boundary in aluminum metal decreases by 95% when the grain boundary energy increases by around 0.35 J/m 50. The intended task of the GANs was to generate these low-interface-energy structures. We assembled 1,500 low-interface-energy structures that were all relaxed by M3GNet + DFT, for use as the GAN training dataset.
2.3 GAN architecture comparison
@@
A typical GAN consists of two neural networks: a generator and a discriminator. The role of the generator is to generate material structures from input noise, whereas the role of the discriminator is to distinguish between the real (training) structures and the fake (generated) structures. The generator and discriminator compete with each other during training to progressively improve the quality of the generated structures. The point cloud representation was used for the GANs, and since all of the interface structures share the same lattice parameters, each structure was simply represented by the fractional coordinates of its atoms. We tested a couple of different GAN architectures as summarized in Fig. 3. We will present each architecture in succession and how they compare against each other.
@@
The base GAN model (CryinGAN) was adapted from the Composition-Conditioned Crystal GAN (CCCGAN) presented by Kim et al. 25, which was used to generate Mg-Mn-O ternary materials. One-dimensional (1D) convolutions were used in the discriminator to extract the latent features of structures, an inspiration taken from PointNet 33, a 3D object classification and segmentation network. These 1D convolutions have been similarly implemented in other point-cloud-based crystal generative models such as FTCP-VAE 9 and CubicGAN 10. We simplified and generalized CCCGAN to be used for any periodic system with fixed lattice, composition, and number of atoms. The discriminator of the original CCCGAN relied solely on atomic coordinates and lattice parameters to distinguish between real and fake structures. To further provide explicit bond distance information to the discriminator, we added a second discriminator to the CryinGAN model that has the same architecture, but accepts bond distances as input instead of coordinates.
@@
The CryinGAN model architecture is shown in Fig. 4. Note that each interface structure has a total of 264 atoms. The generator accepts random gaussian noise as input, and produces fractional coordinates of structures as output. Of the two discriminators, one accepts fractional coordinates as input, and the other accepts bond distances (6 nearest-neighbors of each atom) as input. In the first convolutional layer of the discriminators, the fractional coordinates/bond distances are convoluted along each row separately (i.e., separate 1D convolutions for each atom). CryinGAN implements the Wasserstein loss, which was shown to provide more stable training by preventing the vanishing gradients issue of the traditional GAN 51, 52. The discriminator loss function with a gradient penalty term for improved stability 52 is as follows:
@@
where is the discriminator output, and are the inputs for generated and real structures respectively. The distribution is taken over interpolated samples between the distribution of real structures , and the distribution of generated structures . In the code implementation, is obtained by interpolating between the fractional coordinates/bond distances of training structures and generated structures. The interpolation point between any two data points is chosen randomly. The parameter is the gradient penalty coefficient set to 10 similar to past Wasserstein GANs 52, 10, 25. The total discriminator loss, to be minimized, is a weighted sum of the losses from both discriminators
@@
Here, and are the losses from the fractional coordinate discriminator and bond distance discriminator respectively, and is the weight of the bond distance discriminator loss. The total generator loss, to be maximized, is computed similarly according to the following equations:
@@
CryinGAN models were trained on the assembled dataset of low interface energy structures with varying values. For each value, three separate models were trained and the combined results were used for analysis. A visualization of the training process is shown in ESI Movie S1, where CryinGAN progressively learns to generate low interface energy structures over training epochs. Due to noise in the models, the generated structures often have a small fraction of atoms that are too close to each other; in these cases we manually moved the atoms apart before relaxation. As increases, we found that the number of pairs of atoms generated too close together quickly decreases and then levels off around = 0.05-0.1 (see ESI Fig. S2). This observation is an indicator that including the bond distance discriminator is effective for training the generator to create structures with more reasonable atom-atom distances, in terms of minimal spacing. The generated structures were relaxed using M3GNet to evaluate their interface energies for comparison between models. We found that M3GNet relaxation alone provided sufficiently accurate interface energies without further DFT relaxation (see Table 3 in Methods section).
@@
The objective was to obtain a model with the lowest interface energy distribution. Fig. 5 shows the interface energy distribution for the models trained with different parameters . For each value of , three separate models were trained and the spread of the interface energy distributions is indicated by the shading. The interface energy distribution of all trained models (without shading) is provided in ESI Fig. S3. The results indicate that as increases from 0 to 0.05, the interface energy distribution shifts to lower energies. However, as further increases to 0.1, the interface energy distribution shifts to higher energies. This observation shows that optimal values of improve model performance, but excessive weight on the bond distance discriminator causes the generator to prioritize the bond distances too much over positioning the atoms correctly. The use of a second discriminator does slow down training compared to using only a single discriminator, requiring around twice as long to train the model for the same number of epochs. However, we found that two discriminators still outperform a single discriminator given the same amount of training time (see ESI Fig. S4), justifying the benefits of the bond distance discriminator. Also, whereas a large percentage of structures (around 34%) generated using the models trained with failed to converge in the M3GNet relaxations, only around 11% of structures generated using the models trained with did not converge. These results show that with an appropriately tuned , the bond distance discriminator helps to improve the model performance.
We considered a different architecture that circumvents the need to tune , where both fractional coordinate and bond distance discriminators were combined into a single discriminator, which we refer to as CryinGAN-comb (see ESI Fig. S5). The outputs were combined after the pooling layer, and the model was allowed to learn the relative importance of the coordinate and bond distance latent features on its own through fully connected layers. For CryinGAN-comb, the bond distances were directly obtained using fractional coordinates (instead of Cartesian coordinates), to provide a more direct link to the atomic coordinates which are also represented in fractional coordinates. CryinGAN-comb was trained using the same dataset, then structures were generated from it and relaxed using M3GNet. Compared to CryinGAN trained with , CryinGAN-comb generates structures with fewer atoms too close together (around 2 times fewer pairs of atoms 1.5 Å). However, as shown in Fig. 6a, the interface energy distribution of (relaxed) structures generated from CryinGAN-comb is significantly higher than CryinGAN. For reference, the interface energy distribution of all trained models (without shading) is provided in ESI Fig. S6. These results show that combining the discriminators in CryinGAN-comb leads to less useful discriminator gradients to train the generator, resulting in poor quality of structures being generated. In contrast, separating the discriminators as in CryinGAN helps the GAN to learn more effectively.
Next, we considered different choices of the pooling layer for the discriminator. Wang et al. 53 found that the type of pooling operation affected the sampling sensitivity of the discriminator and the overall performance of the GAN. The sampling sensitivity describes how sensitive the discriminator is to changes in point density or the sampling pattern of the input point cloud. Their results suggest that max pooling produces a discriminator with lower sampling sensitivity than average pooling. The CryinGAN discriminators use average pooling, and we tested an architecture which uses max pooling instead, referred to as CryinGAN-max (see ESI Fig. S7a). We also tested an architecture named CryinGAN-mix that uses the mix pooling operation proposed by Wang et al. 53, where both max and average pooling operations are used together (see ESI Fig. S7b). CryinGAN-max and CryinGAN-mix were trained using = 0 and 0.05. For = 0.05, only the pooling layer of the bond distance discriminator was changed, and we kept the average pooling for the fractional coordinate discriminator. This choice allows us to study the effect of the pooling choice on the two discriminators independently.
The interface energy distributions of relaxed structures generated using CryinGAN, CryinGAN-max, and CryinGAN-mix are shown in Fig. 6b and c. The interface energy distribution of all trained models (without shading) is provided in ESI Fig. S8 for reference. For = 0 (Fig. 6b), we observe that CryinGAN (average pooling) significantly outperforms CryinGAN-max and CryinGAN-mix. Although the use of max pooling was beneficial for the generation of 3D objects where the object shape is the most important aspect to capture 53, it appears that a higher sampling sensitivity is needed for atomic configurations, for which the local coordination environment is the important aspect to capture. For = 0.05 (Fig. 6c), we observe that the distributions are more similar across the three architectures, but CryinGAN still has the lowest energy distribution. The pooling choice of the bond distance discriminator does not relate to the sampling sensitivity like the fractional coordinate discriminator, and the choice appears to have a smaller effect on model performance. Overall, we find that average pooling works best for both discriminators.
@@
Finally, we attempted a graph convolutional discriminator by adapting the commonly used Crystal Graph Convolutional Neural Networks (CGCNN) 54. A higher sampling sensitivity can be achieved through the use of graph convolutions, where each atom is convoluted with its surrounding atoms and bonds. Unlike the PointNet-based 1D convolutional architecture, graph convolutions include explicit information of the neighbor relationship between atoms. However, the training losses diverged and it was not possible to train a GAN that could generate useful structures (see ESI Fig. S9). This result is consistent with the findings of Wang et al. 53, whose graph convolutional GAN also failed to produce point clouds of 3D objects. Graph convolutions are highly sensitive to sampling, making them prone to overfocus on the sampling pattern of a point cloud instead of the overall structure. This overfocus bears similarity to the behavior observed here for CryinGAN at high values. For graph convolutions to be implemented in point-cloud-based GANs, we expect that a carefully designed architecture will be required to take advantage of its sampling sensitivity without destabilizing training.
2.4 Evaluation of CryinGAN-generated structures

### v2

@@ 2.1.3 Li3ScCl6(100)-LiCoO2(110) battery interface
To perform relaxations faster, we trained from scratch a machine learning interatomic potential, M3GNet 49, for the LSC-LCO interfaces. A total of 15,484 training structures consisting of optimized structures and intermediate ionic steps of the DFT relaxations were used to train the M3GNet interatomic potential. The M3GNet model achieved low test set mean absolute errors (MAEs) of 2.70 mev/atom, 20.9 meV/Å, and 0.0146 GPa for energy, force, and stress respectively. Machine learning interatomic potentials with similar (or higher) MAEs showed good performance in relaxations and molecular dynamics simulations when applied to other Li-ion conductors 49, 50. We then relaxed randomly generated interface structures using the M3GNet interatomic potential. The M3GNet-relaxed structures were found to be near DFT convergence (refer to Table 5 in the Methods section). The interface energies of the relaxed structures were distributed across a wide range (approximately 1.4 J/m2) as shown in ESI Fig. S4. We define structures with interface energies no higher than 0.4 J/m2 relative to the lowest energy structure to be low-interface-energy structures. For reference, the observation frequency of a grain boundary in aluminum metal decreases by 95 % when the grain boundary energy increases by around 0.35 J/m2 51. We assembled 1,500 low-interface-energy structures (all relaxed by M3GNet only) as the disordered interface dataset.
2.2 Generative models
Five generative models were benchmarked on Dismai-Bench. Four of these (CDVAE 26, DiffCSP 27, CrysTens 28, UniMat 29) are existing models, and one (CryinGAN) was developed as part of this work to demonstrate the application of Dismai-Bench in model development. An overview of the first four models is outlined here, whereas CryinGAN will be presented in detail in Section 2.4.
CDVAE 26 is a VAE that uses equivariant graph neural networks for its encoder and decoder. The encoder is a graph convolutional network that encodes material structures into latent representations (). Three multilayer perceptrons (MLPs) are used to predict the composition, lattice parameters, and number of atoms from . These predictions are used to initialize structures (corresponding to the sampled ), where the atoms are initialized at random positions. The decoder is a graph diffusion model that denoises both the atomic coordinates and atomic species of the atoms.
In our early tests of training CDVAE on Dismai-Bench datasets, CDVAE was found to fail in generating structures with the correct compositions (see ESI Fig. S5). Although CDVAE was able to predict the compositions correctly from , the compositions became incorrect after the atomic species of the atoms were denoised. Therefore, we modified CDVAE such that the atomic species denoising becomes an optional feature, and all CDVAE benchmarking was performed without atomic species denoising. We also tested the effect of atomic species denoising when CDVAE is trained on the MP-20 dataset 26, which includes structures from the Materials Project 52 of various compositions. Here, denoising the atomic species was found to increase the composition accuracy of reconstructed structures from around 24 % to 54 %. However, denoising the atomic species appears to be detrimental for larger structures (such as those in the Dismai-Bench datasets), and fails even when all structures have the same composition.
DiffCSP 27 is another graph diffusion model. The main feature introduced in DiffCSP is the ability to jointly denoise the atomic coordinates and lattice parameters. In contrast, CDVAE predicts the lattice parameters first, and they remain fixed throughout the diffusion steps. Note that DiffCSP does not denoise the atomic species. To better capture periodicity, DiffCSP also uses fractional instead of Cartesian coordinates (as in CDVAE), and uses periodic translation invariant Fourier transformations in its message passing. However, DiffCSP does not include bond angle information in its graphs, whereas CDVAE does.
When DiffCSP was trained on the disordered interface dataset allowing joint atomic coordinate and lattice parameter diffusion, the generated structures were found to be of poor quality (refer to ESI Fig. S6a). We modified DiffCSP to allow teacher forcing of lattice parameters during the initial training epochs, where the ground truth lattice parameters are used as input, and lattice cost is not used to update the model (only coordinate cost is used). The quality of the generated structures improved when trained with teacher forcing (see ESI Fig. S6b). However, structures with the best quality were still obtained when DiffCSP was trained without any lattice diffusion, using the ground truth lattice parameters as input (see ESI Fig. S6c). Therefore, all benchmarking of DiffCSP presented in Section 2.3 was performed without lattice denoising. This choice also provides a consistent comparison between DiffCSP and CDVAE, since CDVAE also does not perform lattice denoising. For reference, the Dismai-Bench metrics of DiffCSP trained with lattice denoising and teacher forcing are listed in ESI Table S1. DiffCSP performance drops with lattice denoising across the metrics (see Section 2.3 for benchmarking details).
CrysTens 28 is an image-like representation for materials. The pixel values of a CrysTens image are filled with information of the structure such as lattice parameters, fractional coordinates, and atomic number. Each CrysTens image has four channels, analogous to the RGB color channels of an image. The first channel includes a pairwise distance matrix between all atom pairs, and the remaining three channels include the pairwise , , and matrices respectively. CrysTens is used with a 2D U-Net diffusion model for image generation, and the generated CrysTens images are reconstructed back into material structures. We made no major modifications to the baseline implementation. The dimensions of the CrysTens images were increased to fit Dismai-Bench structures (original implementation only allowed up to 52 atoms). Refer to the Methods section for more details.
UniMat 29 is a video-like representation for materials. Each frame of a UniMat video represents a single atom in the structure, and has three channels corresponding to the , , and coordinates. The atomic species of the atom is indicated by the pixel location in the frame (e.g., top left pixel is H), where those pixel values correspond to the coordinates of the atom, and all other pixel values are set to -1. UniMat is used with a 3D U-Net diffusion model for video generation, and the generated UniMat videos are reconstructed back into materials. Despite the lack of any geometrical information beyond atomic coordinates, UniMat was shown to outperform CDVAE in discovering materials with lower formation energy (upon DFT relaxation of the generated structures). As the UniMat code is currently not openly available, we used an open-access implementation 53 of the 3D U-Net model 54 that the UniMat model was repurposed from. Besides this change, no major modifications were made to UniMat. The dimensions of the UniMat frames were decreased, since each Dismai-Bench structure has only five atomic species at most.
2.3 Dismai-Bench
Benchmarking was performed by training all generative models on each dataset separately from scratch, such that the models only generate one type of structure at a time. 1,000 structures were generated using each model, and post-processed as described in the Methods section. The disordered interface structures were relaxed using the M3GNet 49 interatomic potential, and the amorphous Si structures were relaxed using the SOAP-GAP 39 interatomic potential. No relaxations were performed on the alloy structures. These final structures were used to calculate the benchmark metrics. For each generative model architecture, three separate models were trained for each dataset, and the metrics were averaged across the three models.
Although the benchmark metrics for all five generative model architectures are listed in this section, only the results for CDVAE, DiffCSP, CrysTens, and UniMat will be discussed here, since the CryinGAN architecture has not been presented yet (see Section 2.4). The benchmark results for CryinGAN will be discussed in Section 2.6, along with an overall comparison of the general performance of all architectures.
2.3.1 Disordered LSC-LCO interface
Examples of interface structures generated by the models are shown in Fig. 2. The structures shown are as-generated without any post-processing or relaxation. Visually, the structures generated by CryinGAN, CDVAE, and DiffCSP are similar to the training structures. The coordinate-based U-Net diffusion models generated the most noisy structures.
| Model |
(min, max) |
(min, max) |
(min, max) |
(min, max) |
% struc failed
(min, max) |
| CDVAE |
0.0664
(0.0601, 0.0697) |
0.0427
(0.0372, 0.0504) |
0.201
(0.186, 0.214) |
0.0547
(0.0508, 0.0589) |
0.00
(0.00, 0.00) |
| DiffCSP |
0.0439
(0.0414, 0.0474) |
0.0272
(0.0240, 0.0298) |
0.0954
(0.0909, 0.102) |
0.0370
(0.0353, 0.0388) |
7.17
(6.00, 8.90) |
| CrysTens |
0.0202
(0.0160, 0.0244) |
0.0141
(0.00707, 0.0239) |
0.0972
(0.0888, 0.106) |
0.0213
(0.0164, 0.0257) |
55.3
(50.6, 61.7) |
| UniMat |
0.131
(0.0966, 0.154) |
0.235
(0.165, 0.292) |
0.101
(0.0891, 0.111) |
0.151
(0.111, 0.181) |
64.6
(60.6, 68.3) |
| CryinGAN |
0.0538
(0.0466, 0.0600) |
0.0274
(0.0210, 0.0359) |
0.0888
(0.0838, 0.0963) |
0.0426
(0.0379, 0.0451) |
9.00
(7.20, 10.2) |
The generated structures were relaxed, and the benchmark metrics were calculated by analyzing the local coordination environment (motifs) of the atoms. The CrystalNNFingerprint 55, which contains the coordination likelihoods and local structure order parameters of a given atom, was averaged across the structures generated by each model, and separately obtained for the training dataset as well. We calculated the fingerprints of the cations (Li, Co, and Sc) coordinated to the anions (Cl and O). We also appended the fraction of Cl and O neighbors in each motif to the fingerprints, so that the fingerprints contain both chemical and coordination information. For each model, the Euclidean distance between the average fingerprint of the training structures and the generated structures was calculated.
The benchmark metrics for the disordered interfaces are shown in Table 1. Parameters , , , and represent the fingerprint Euclidean distance of Li, Co, Sc, and all cations respectively. The percentages of structures that failed to be post-processed or relaxed are also listed. Notably, CDVAE is the only model that generates structures close to being fully relaxed, requiring only an average of 10 relaxation steps per structure (see ESI Table S2). All other models required at least 68 relaxation steps on average, and the U-Net models have % failed structures. UniMat exhibited the highest % failed structures and , likely because it uses a less expressive material representation (atomic coordinates only). CrysTens, despite the inclusion of bond distances, also has high % failed structures. Although its Euclidean distance metrics are small, the low values are only achieved after relaxation (refer to Fig. 2 and Table S2 for unrelaxed structures). On the other hand, the graph diffusion models perform significantly better than the U-Net diffusion models due to the higher expressive power of graphs. Comparing between CDVAE and DiffCSP, CDVAE generates interface structures closer to convergence than DiffCSP, but upon relaxation, DiffCSP achieves lower distance metrics than CDVAE.
2.3.2 Amorphous silicon
Examples of amorphous Si structures generated by the models are shown in Fig. 3. Only the graph diffusion models (CDVAE & DiffCSP) were able to generate the structures successfully. The other models generated random-looking structures with large voids and little bonding between atoms. Therefore, further detailed benchmarking of amorphous Si was only performed for CDVAE and DiffCSP. Similar to the disordered interfaces, we determined the coordination motif fingerprints of Si, and the Euclidean distance between the average fingerprint of the training and generated structures. We also determined the radial distribution functions (RDFs) and bond angle distributions of the structures, and compared them to that of the training set.
| Model |
(min, max) |
(min, max) |
(min, max) |
% struc failed
(min, max) |
| CDVAE |
0.0402
(0.0396, 0.0412) |
0.392
(0.381, 0.408) |
0.00332
(0.00312, 0.00353) |
0.00
(0.00, 0.00) |
| DiffCSP |
0.0647
(0.0462, 0.0908) |
1.39
(1.03, 1.69) |
0.0103
(0.00797, 0.0133) |
0.00
(0.00, 0.00) |
The benchmark metrics for amorphous Si are shown in Table 2. Parameters , , and represent the Euclidean distance for motif fingerprint, RDF, and bond angle distribution respectively. Both CDVAE and DiffCSP had 0 % failed structures, but CDVAE-generated structures were significantly closer to being fully relaxed than DiffCSP-generated structures. CDVAE only required an average of 23 relaxation steps per structure, whereas DiffCSP required an average of 140 steps per structure (see ESI Table S3). Similarly, all distance metrics of CDVAE are lower than DiffCSP. The lack of bond angle information in DiffCSP’s graphs is likely one main contributor to these trends. The RDFs and bond angle distributions of CDVAE and DiffCSP are shown in ESI Fig. S7. DiffCSP and CDVAE encode distance information in their graphs, and both show similar RDFs to the training dataset. However, there is a clear difference when comparing the bond angle distributions, where DiffCSP shows a larger discrepancy with the training dataset than CDVAE. These results indicate that bond angle information is particularly beneficial for helping generative models learn amorphous structures, since amorphous structures have more complicated arrangements of atoms that do not locate on lattice sites.
2.3.3 Disordered stainless steel alloy
Examples of disordered alloy structures generated by the models are shown in Fig. 4. CryinGAN, CDVAE, and DiffCSP were able to generate the FCC alloy structures successfully, whereas CrysTens and UniMat were unable to reproduce the underlying FCC lattice. Therefore, detailed alloy benchmarking was only performed for CryinGAN, CDVAE, and DiffCSP. Interestingly, CDVAE was prone to generating rotated structures (since graphs are rotationally invariant), but all DiffCSP-generated structures were unrotated. Some CDVAE-generated structures had noisy lattices, where atoms exhibited relatively large deviations from pristine FCC lattice sites (see ESI Fig. S8a), whereas all DiffCSP-generated structures had well-defined lattices. A fraction of CDVAE-generated structures had slightly shorter lattice spacing, creating additional sites that were vacant since the total number of atoms was fixed (see ESI Fig. S8b). DiffCSP’s stronger ability to learn the FCC lattice is likely due to the use of fractional coordinates and Fourier transformations in its message passing to capture periodicity 27.
| Model |
(min, max) |
% struc w/ vac
(min, max) |
% struc failed
(min, max) |
| 300 K, narrow SRO | |||
| CDVAE |
0.0604
(0.0540, 0.0675) |
35.8
(32.0, 41.2) |
0.13
(0.00, 0.40) |
| DiffCSP |
0.0645
(0.0566, 0.0792) |
94.3
(91.8, 96.5) |
0.00
(0.00, 0.00) |
| CryinGAN |
0.117
(0.106, 0.126) |
100
(100, 100) |
0.00
(0.00, 0.00) |
| 300 K, wide SRO | |||
| CDVAE |
0.0658
(0.0598, 0.0742) |
37.4
(34.0, 43.4) |
0.07
(0.00, 0.10) |
| DiffCSP |
0.103
(0.0831, 0.121) |
95.8
(95.4, 96.2) |
0.00
(0.00, 0.00) |
| CryinGAN |
0.125
(0.119, 0.129) |
100
(100, 100) |
0.00
(0.00, 0.00) |
| 1500 K, narrow SRO | |||
| CDVAE |
0.0621
(0.0589, 0.0665) |
54.8
(51.9, 56.8) |
0.37
(0.20, 0.50) |
| DiffCSP |
0.0308
(0.0304, 0.0313) |
91.7
(91.5, 91.8) |
0.00
(0.00, 0.00) |
| CryinGAN |
0.0643
(0.0629, 0.0659) |
100
(100, 100) |
0.00
(0.00, 0.00) |
| 1500 K, wide SRO | |||
| CDVAE |
0.0618
(0.0549, 0.0694) |
61.3
(55.9 68.2) |
0.27
(0.20, 0.30) |
| DiffCSP |
0.0338
(0.0332, 0.0342) |
92.3
(89.2, 94.4) |
0.00
(0.00, 0.00) |
| CryinGAN |
0.0650
(0.0609, 0.0691) |
100
(100, 100) |
0.00
(0.00, 0.00) |
The generated structures were post-processed to remove atoms not on lattice sites. Any structure with > 50 atoms removed (20 % of all atoms) was considered a failed structure and rejected. Then, the fingerprint of each structure was calculated using a vector of conditional probabilities of observing each cluster (monomer/dimer) in the structure, up to the 7 neighbor shell. The Euclidean distance between the average cluster fingerprint of the training and generated structures, , was calculated for each model.
Benchmark metrics for the disordered alloys are shown in Table 3. The percentage of failed structures is 0 % for DiffCSP and CryinGAN, while CDVAE exhibits a tiny percentage of failed structures. However, the percentage of structures with site vacancies is 90-100 % for DiffCSP and CryinGAN. In comparison, CDVAE has significantly lower percentages of structures with vacancies, around roughly 40 % for 300 K structures and 60 % for 1500 K structures. Although DiffCSP-generated structures had more well-defined lattices than CDVAE-generated structures, almost all structures had overlapping atoms on lattice sites, resulting in site vacancies in the structures.
For the 300 K structures with narrow SRO, the values of CDVAE and DiffCSP are similar. However, for the 300 K structures with wide SRO, CDVAE has lower than DiffCSP. The wide SRO structures have less consistent SRO distributions, so DiffCSP had more difficulty in learning the wide SRO structures than the narrow SRO structures. On the other hand, only increased slightly for CDVAE between the wide and narrow SRO structures, indicating CDVAE’s stronger ability to learn different degrees of SRO. For the 1500 K structures, which more resemble random solid solutions, there is little difference in between the narrow and wide SRO structures. Here, is lower for DiffCSP than CDVAE. Some factors contributing to DiffCSP’s better performance in generating random solid solutions may be its stronger ability in learning the FCC lattice, weaker ability in learning SRO patterns, and low number of site vacancies (1.5 vacancies per structure). The Warren-Cowley SRO parameter distributions for the 1st and 2nd nearest neighbor interactions are shown in ESI Fig. S9-11 for reference. Overall, the models were able to generate structures with similar SRO distributions to the training structures.
2.4 CryinGAN development
We present here a case study of developing a generative model with the help of Dismai-Bench, demonstrating how meaningful feedback about model performance is obtained through direct comparisons between generated and training structures. We considered the case of a point-cloud-based GAN to generate low-interface-energy LSC-LCO interface structures. We chose to focus on the interface structures during development for simplicity, and subsequently benchmarked CryinGAN on other datasets to evaluate its generalizability. The disordered interfaces were also the only structures where all four diffusion models were able to generate successfully, hence providing the most informative model comparison. We show that, given the correct architecture, a coordinate-based representation can still perform well on Dismai-Bench (unlike CrysTens and UniMat), and that more complicated architectures do not necessarily outperform simpler architectures.
@@ 2.4 CryinGAN development
A typical GAN consists of two neural networks: a generator and a discriminator. The role of the generator is to generate material structures from input noise, whereas the role of the discriminator is to distinguish between the real (training) structures and the fake (generated) structures. The generator and discriminator compete with each other during training to progressively improve the quality of the generated structures. The point cloud representation was used for the GANs, and since all of the structures in the dataset share the same lattice parameters, each structure was represented by the fractional coordinates of its atoms only. We tested a couple of different GAN architectures as summarized in Fig. 5.
@@
The base GAN model (CryinGAN) was adapted from the Composition-Conditioned Crystal GAN (CCCGAN) presented by Kim et al. 25, which was used to generate Mg-Mn-O ternary materials. One-dimensional (1D) convolutions were used in the discriminator to extract the latent features of structures, an inspiration taken from PointNet 34, a 3D object classification and segmentation network. These 1D convolutions have been similarly implemented in other point-cloud-based crystal generative models such as FTCP-VAE 9 and CubicGAN 10. We simplified and generalized CCCGAN to be used for any periodic system with fixed lattice, composition, and number of atoms. We did not include any conditional generation capability in CryinGAN for simplicity. The discriminator of the original CCCGAN relied solely on atomic coordinates and lattice parameters to distinguish between real and fake structures. To further provide explicit bond distance information to the discriminator, we added a second discriminator to the CryinGAN model that has the same architecture, but accepts bond distances as input instead of coordinates.
@@
The CryinGAN model architecture is shown in Fig. 6. The generator accepts random gaussian noise as input, and produces fractional coordinates of structures as output. Of the two discriminators, one accepts fractional coordinates as input, and the other accepts bond distances (6 nearest-neighbors of each atom) as input. In the first convolutional layer of the discriminators, the fractional coordinates/bond distances are convoluted along each row separately (i.e., separate 1D convolutions for each atom). Note that CryinGAN is permutationally invariant to atom ordering within each atomic species block (e.g., order of the 72 Li atoms does not matter). CryinGAN implements the Wasserstein loss, which was shown to provide more stable training by preventing the vanishing gradients of the traditional GAN 56, 57. The discriminator loss function with a gradient penalty term for improved stability 57 is:
@@
where is the discriminator output, and are the inputs for generated and real structures respectively. The distribution is taken over interpolated samples between the distribution of real structures , and the distribution of generated structures . In the code implementation, is obtained by interpolating between the fractional coordinates/bond distances of training structures and generated structures. The interpolation point between any two data points is chosen randomly. The parameter is the gradient penalty coefficient set to 10 similar to past Wasserstein GANs 57, 10, 25. The total discriminator loss, to be minimized, is a weighted sum of the losses from both discriminators
@@
Here, and are the losses from the fractional coordinate discriminator and bond distance discriminator respectively, and is the weight of the bond distance discriminator loss. The total generator loss, to be maximized, is computed similarly according to:
@@
We developed the GAN models with the intention of evaluating the best performing model by comparing between DFT-relaxed training and generated structures. The training dataset was curated from structures relaxed with the M3GNet interatomic potential followed by DFT calculations (refer to Methods section). Note that the CryinGAN Dismai-Bench metrics shown in Section 2.3 were calculated using the same procedure as described previously. To study the effect of on CryinGAN model performance, we trained CryinGAN models on the (DFT-relaxed) dataset of interface structures with varying . A visualization of the training process is shown in ESI Movie S1, where CryinGAN progressively learns to generate low interface energy structures over training epochs. Due to noise in the models, the generated structures often have a small fraction of atoms that are too close to each other. As increases, we found that the number of pairs of atoms generated too close together quickly decreases and then levels off around = 0.05-0.1 (see ESI Fig. S12). This observation is an indicator that including the bond distance discriminator is helpful for training the generator to create structures with more reasonable atom-atom distances. The generated structures were then relaxed using M3GNet, refer to ESI Fig. S13 for examples of structures before and after relaxation.
@@
The objective was to obtain a model with the lowest interface energy distribution. Fig. 7 shows the interface energy distribution for the models trained with different . For each , three separate models were trained and the spread of the interface energy distributions is indicated by the shading (the distribution of all trained models without shading is provided in ESI Fig. S14). The results indicate that as increases from 0 to 0.05, the interface energy distribution shifts to lower energies. However, as further increases to 0.1, the interface energy distribution shifts to higher energies. The coordination motif fingerprint distance between the generated and training structures also shows the same trend (see ESI Fig. S15), where it decreases between and , then increases when . These observations show that optimal values of improve model performance, but excessive weight on the bond distance discriminator causes the generator to prioritize the bond distances too much over positioning the atoms correctly. The use of a second discriminator does slow down training compared to using only a single discriminator, requiring around twice as long to train the model for the same number of epochs. However, we found that two discriminators still outperform a single discriminator given the same amount of training time (see ESI Fig. S16), justifying the benefits of the bond distance discriminator. Whereas around 34 % of the generated structures failed to converge during M3GNet relaxation for , only around 11 % did not converge for . These results show that with an appropriately tuned , the bond distance discriminator improves model performance.
We further considered a couple of alternative GAN architectures (refer to Fig. 5). CryinGAN-comb combines both discriminators into a single discriminator, circumventing the need to tune . CryinGAN-max uses max pooling in the discriminator, instead of average pooling as in CryinGAN. CryinGAN-mix uses the mix pooling operation proposed by Wang et al. 58, where both max and average pooling operations are used together. The type of pooling operation affects the sampling sensitivity of the discriminator and the overall performance of the GAN. The sampling sensitivity describes how sensitive the discriminator is to changes in point density or the sampling pattern of the input point cloud. Max pooling was reported to cause lower sampling sensitivity than average pooling 58. Overall, CryinGAN was found to outperform all of these alternative architectures, see ESI Supplementary Note 2 for further details.
@@
The superior performance of CryinGAN over CryinGAN-max and CryinGAN-mix shows that higher sampling sensitivity is beneficial for learning atomic configurations. The sampling sensitivity can be further increased through the use of graph convolutions, where each atom is convoluted with its surrounding atoms and bonds. We attempted a graph convolutional discriminator by adapting the commonly used Crystal Graph Convolutional Neural Networks (CGCNN) 59. However, the training losses diverged and it was not possible to train a GAN that could generate useful structures (see ESI Fig. S17). This result is consistent with the findings of Wang et al. 58, whose graph convolutional GAN also failed to produce point clouds of 3D objects. Graph convolutions are highly sensitive to sampling, making them prone to overfocus on the sampling pattern of a point cloud instead of the overall structure. This overfocus bears similarity to the behavior observed here for CryinGAN at high values. For graph convolutions to be implemented in point-cloud-based GANs, we expect that a carefully designed architecture will be required to take advantage of its sampling sensitivity without destabilizing training.
2.5 Detailed evaluation of CryinGAN-generated interfaces